OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)).

OpenStudy (lgbasallote):

first what is g(2)?

OpenStudy (anonymous):

???

OpenStudy (lgbasallote):

what do you mean "????"

OpenStudy (anonymous):

your question said what is g(2) that how the poblem was given

OpenStudy (anonymous):

First you must find g(2) by putting into the equation for g(2). Then you would take the step as to put what g(2) = into the f(x) equation.

OpenStudy (anonymous):

By putting 2 into the equation for g, excuse me. Its a multi-step equation.

OpenStudy (anonymous):

When you see g(2), that means solve g(x) where x = 2. \[\Large g(x) = x^{2}-6x-7\] \[\Large g(2) = 2^{2}-(6*2)-7= ?\]

OpenStudy (lgbasallote):

f(g(2) ) means you find g(2) then substitute that to x in f(g(2) )

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Note: Alternatively you could solve it all at once (but that may confusing right off the bat): \[\Large f(g(2)) = (x^{2}-6x-7)+8 = x^{2}-6x+1\]

OpenStudy (anonymous):

What did you get for g(2)?

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

Correct. Since g(2)=-15, that means f(g(2)) = f(-15). \[\Large f(x) = x+8\] \[\Large f(-15) = -15+8 = ?\]

OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). first find g(2) then once you find that plug that to f(x)